OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2.

OpenStudy (texaschic101):

slope intercept form is y = mx + b y - y1 = m(x - x1) y - 2 = 1/2(x - 4) y - 2 = 1/2x - 2 y = 1/2x - 2 + 2 y = 1/2x + 0 or just y = 1/2x